OpenStudy (anonymous):

-6-9x=-9x-4 is my answer 0 right

jimthompson5910 (jim_thompson5910):

what happens when you add 9x to both sides?

OpenStudy (anonymous):

0

jimthompson5910 (jim_thompson5910):

the -9x on the left side goes away, so does the 9x on the right

jimthompson5910 (jim_thompson5910):

leaving you with -6 = -4

OpenStudy (anonymous):

what bout variable

jimthompson5910 (jim_thompson5910):

what do you mean?

OpenStudy (anonymous):

your answer has no vafriable

jimthompson5910 (jim_thompson5910):

is -6 = -4 true? or false?

OpenStudy (anonymous):

let me plug in for x and check one mine

OpenStudy (anonymous):

what that means ? @jim_thompson5910

OpenStudy (anonymous):

variable supposed to be on one side right

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

say you had something simple like x = x that's always true

jimthompson5910 (jim_thompson5910):

x = x+1 is false because there is no way you can have a number equal the next number after that same number (eg: there's no way 5 could equal 6)

OpenStudy (anonymous):

But if you want to know the answer, just plug in the number you got (Ex: For you, 0) for the variables

jimthompson5910 (jim_thompson5910):

x = x+1 x-x = x+1-x ... Subtract x from both sides. 0 = 1 ... this is false, so there are no solutions

OpenStudy (anonymous):

ok thx guys